# Dimitrios M. Thilikos ${}^{,}$

### v3

@@ 5.1 A local structure theorem for excluding Dyck-grids
Let us first assume that $\mathbb{S}=\{\Sigma^{\varnothing}\}.$ In this case $\mathsf{sobs}(\mathbb{S})$ contains only surfaces homeomorphic to the sphere. Thus, the orientable Dyck-grid and the non-orientable one are both of Euler-genus zero. Therefore, this case follows immediately from Subsection 3.1.
From now on we may assume $\mathbb{S}$ to contain a non-void surface.
Suppose $g_{2}=0.$ This means that $\mathsf{sobs}(\mathbb{S})$ contains only orientable surfaces of Euler-genus $g_{1}.$ Notice that this implies that $\mathbb{S}$ contains a non-orientable surface of Euler-genus $g_{1}$ and $g_{1}+1$.
Let $g_{1}=2h_{1}.$ Now let $k\in\mathbb{N}$ be some integer and $G$ be some graph that excludes $\mathscr{D}^{(h_{1},0)}_{k}$ as a minor. In case $k\in[0,1]$ this graph is planar, and thus our claim follows from Subsection 3.1. Thus, we may assume $k\geq 2.$ Finally, let $t\coloneqq 12(162^{2g_{1}}+4)k.$ Let
$\mathbf{a}_{g_{1},0}$ $\coloneqq t^{10^{7}t^{26}}$
$\mathbf{g}_{\mathsf{o},g_{1},0}$ $\coloneqq g_{1}-1$
$\mathbf{g}_{\mathsf{n},g_{1},0}$ $\coloneqq{g_{1}}$
$\mathbf{v}_{\mathsf{b},{g_{1},0}}$ $\coloneqq 2t^{2}$
$\mathbf{v}_{\mathsf{d},{g_{1},0}}$ $\coloneqq t^{10^{7}t^{26}}$
$\mathbf{R}_{g_{1},0}$ $\coloneqq 49152t^{24}(r+16t)+\left\lceil\frac{5}{2}t^{10^{7}t^{26}}\right\rceil.$
Moreover, let $\mathfrak{D}\coloneqq\{\mathscr{D}^{(h_{1},0)}=\langle\mathscr{D}^{(h_{1},0)}_{i}\rangle_{i\in\mathbb{N}},\}.$ Then $\mathfrak{S}(\mathfrak{D}_{g_{1},0},\mathbf{a}_{g_{1},0},\mathbf{g}_{\mathsf{o},g_{1},0},\mathbf{g}_{\mathsf{n},g_{1},0},\mathbf{v}_{\mathsf{b},{g_{1},0}},\mathbf{v}_{\mathsf{d},{g_{1},0}},\mathbf{R}_{g_{1},0})$ is valid by Theorem 5.1.
Our claim now follows immediately from Theorem 4.2.
Hence, we may now assume $\mathsf{sobs}(\mathbb{S})$ to contain a non-void surface which is non-orientable and one that is orientable. Thus, there exist integers $g_{1}$ and $g_{2}$ where $g_{2}\neq 0$ such that every orientable surface in $\mathsf{sobs}(\mathbb{S})$ has Euler-genus $g_{1}$ and every non-orientable surface in $\mathsf{sobs}(\mathbb{S})$ has Euler-genus $g_{2}.$ In particular, we may assume for every orientable $\Sigma\in\mathsf{sobs}(\mathbb{S})$ and every non-orientable $\Sigma^{\prime}\in\mathsf{sobs}(\mathbb{S})$ that $\Sigma\not\preceq\Sigma^{\prime}$ and $\Sigma^{\prime}\not\preceq\Sigma.$
Let $g_{1}=2h_{1}$ and $g_{2}=2h_{2}+c_{2}$ where $c_{2}\in[2].$ From the discussion above it follows that $h_{2}<h_{1}.$ Now let $k\in\mathbb{N}$ be some integer and $G$ be some graph that excludes both $\mathscr{D}^{(h_{1},0)}_{k}$ and $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as minors. In case $k\in[0,1]$ both of these graphs are planar, and thus our claim follows from Subsection 3.1. Thus, we may assume $k\geq 2.$ Finally, let $g\coloneqq\max\{g_{1},g_{2}\}$ and $t\coloneqq 12(162^{2g}+4)k.$ Let $\mathbf{a}_{g_{1},g_{2}}\coloneqq\mathbf{a}_{g_{1},0}$, $\mathbf{g}_{\mathsf{o},g_{1},g_{2}}\coloneqq\mathbf{g}_{\mathsf{o},g_{1},0}$, $\mathbf{v}_{\mathsf{b},{g_{1},g_{2}}}\coloneqq\mathbf{v}_{\mathsf{b},{g_{1},0}}$, $\mathbf{v}_{\mathsf{d},{g_{1},g_{2}}}\coloneqq\mathbf{v}_{\mathsf{d},{g_{1},0}}$, $\mathbf{R}_{g_{1},g_{2}}\coloneqq\mathbf{R}_{g_{1},0}$, and
$\mathbf{g}_{\mathsf{n},g_{1},g_{2}}\coloneqq{g_{2}-1}$
Moreover, let $\mathfrak{D}\coloneqq\{\mathscr{D}^{(h_{1},0)}=\langle\mathscr{D}^{(h_{1},0)}_{i}\rangle_{i\in\mathbb{N}},\mathscr{D}^{(h_{2},c_{2})}=\langle\mathscr{D}^{(h_{2},c_{2})}_{i}\rangle_{i\in\mathbb{N}}\}.$ Then, by Theorem 5.1, the local structure theorem $\mathfrak{S}_{(g_{1},g_{2})}(\mathfrak{D}_{g_{1},g_{2}},\mathbf{a}_{g_{1},g_{2}},\mathbf{g}_{\mathsf{o},g_{1},g_{2}},\mathbf{g}_{\mathsf{n},g_{1},g_{2}},\mathbf{v}_{\mathsf{b},{g_{1},g_{2}}},\mathbf{v}_{\mathsf{d},{g_{1},g_{2}}},\mathbf{R}_{g_{1},g_{2}})$ is valid.
Our claim now follows immediately from Theorem 4.2. ∎
6 Bounding annotated treewidth
Recall that Theorem 5.2 provides a tree-decomposition $(T,\beta)$ where for each $t\in V(T)$ there exists a set $A\subseteq\beta(t)$ such that the torso $G_{t}$ of $G$ at $t$ has a $\Sigma$-decomposition $d$ of width at most $f(k)$ and breadth at most $f(k)$ after deleting $A$. In the statement of the same lemma we also considered the set $X_{t}$ containing all vertices of $G_{t}-A_{t}$ which are drawn in some vortex of $\delta$ together with the vertices of $A_{t}.$ The purpose of this section is to prove that $X_{t}$ is a low-treewidth modulator of the torso $G_{d}$ whose annotated treewidth is bounded by a function the width and the breath of $\delta$ as well as the Euler-genus of $\Sigma.$ This will permit us to substitute the use of the apex vertices $A_{d}$ and the vortices of $\delta$ by a single low-treewidth modulator.
In this section we present several results on the biggest grid parameter ${\sf bg}$ and its annotated version. Some of these results will also be useful for the proofs of the lower bounds in Section 7. We start with the formal definitions of annotated extension and ${\sf bg}.$
Annotated extensions of parameters.
We use $\mathcal{A}_{\text{all}}$ in order to denote the set of all annotated graphs. Given a minor-monotone graph parameter ${\sf p}:\mathcal{G}_{\rm all}\to\mathbb{N},$ we define its annotated extension ${\sf p}:\mathcal{A}_{\rm all}\to\mathbb{N}$ so that
${\sf p}(G,X)\coloneqq\max\{{\sf p}(H)\mid\mbox{$H$ is an $X$-minor of $G$}\}.$
Notice that $0={\sf p}(G,\emptyset)\leq{\sf p}(G,X)\leq{\sf p}(G,V(G))={\sf p}(G).$
In this paper we apply the above definition for ${\sf tw}$ and ${\sf bg}$ as defined below.
The ${\sf bg}$ parameter.
The biggest grid parameter ${\sf bg}$ outputs the maximum $k$ for which $G$ contains a $(k\times k)$-grid as a minor. Recall that $\{\Sigma^{0,0}\}\mbox{-}\mathsf{bg}$ is the maximum $k$ for which $G$ contains the annulus grid $\mathscr{A}_{k}$ as a minor, therefore ${\sf bg}\sim_{\mathsf{L}}\{\Sigma^{0,0}\}\mbox{-}\mathsf{bg}.$
Notice that the treewidth and the biggest grid parameters are asymptotically equivalent, in fact it holds that ${\sf tw}\sim_{{\mathsf{P}}}{\sf bg}$ (by Subsection 3.1 and the asymptotic equivalence between grids and walls). In this section we study the annotated extension of ${\sf bg}$: for every annotated graph $(G,X),$ ${\sf bg}(G,X)$ is the biggest $k$ such that $G$ contains a $(k\times k)$-grid as an $X$-minor.
Notice that ${\sf bg}(G,X)\geq k$ if and only if there is a collection $\mathcal{S}=\{S_{1},\ldots,S_{k^{2}}\}$ of connected and pairwise-disjoint vertex sets of $G,$ each containing at least one vertex in $X$ and such that the contraction, in $G[\boldsymbol{\boldsymbol{\bigcup}}\mathcal{S}],$ of all edges with both endpoints in the same $S_{i}$ results in a graph containing a $(k\times k)$-grid as a spanning subgraph.
6.1 Walls in surfaces and their radii
An annotated $r$-wall is a pair $(W,S)$ where
  • $W$ is an $r$-wall,
  • every path of $W$ between two vertices of degree three in $W$ contains exactly one vertex from $S.$
  • all vertices of $S$ have degree two in $W.$
A hairy annotated $r$-wall is a pair $(\hat{W},X)$ such that $\hat{W}=W\cup\boldsymbol{\boldsymbol{\bigcup}}\mathcal{P}$ and $X\subseteq V(\hat{W})$ where
  • $W$ is an $r$-wall that contains a vertex set $S\subseteq V(W)$ where $|S|=|X|$ and such that $(W,S)$ is an annotated $r$-wall,
  • $\mathcal{P}$ is a collection of $|X|$ vertex-disjoint paths each with one endpoint in $S$ and with the other in $X$ and such that $V(W)\cap\bigcup\mathcal{P}=S.$
The branch vertices of $\hat{W}$ are the branch vertices of the $r$-wall $W.$
The following observation follows easily by the fact that every $r$-wall is a subdivision of a subgraph of a $(2k\times 2k)$-grid.
[…]
There is a universal constant $c$ such that for every graph $G$ and every set $X\subseteq V(G),$ if ${\sf bg}(G,X)\geq c{k},$ then $G$ contains a hairy annotated $k$-wall $(W,X)$ as a subgraph.
Figure 13: A hairy annotated $5$-wall $(\hat{W},X).$ The edges and the internal nodes of the paths in $\mathcal{P}$ are orangle. The set $S$ is depicted by the magenta vertices and the set $X$ is depicted by the blue circles. The branch vertices are the square vertices.
We say that a vertex of a graph is a branch vertex if it has degree at least 3. Given two vertices $x,y,$ we define their branch distance, denoted by $\mathsf{bdist}_{G}(x,y)$ as the minimum number of branch vertices of a path between $x$ and $y.$ The branch radius of $G$ is at most $r$ if there is an $x\in V(G)$ such that for every $y\in V(G),$ $\mathsf{bdist}_{G}(x,y)\leq r.$
[…]
Let $(\hat{W},X)$ be a hairy annotated $r$-wall and let $Q$ be the graph obtained by adding a $v_{\text{new}}$ and making it adjacent to all vertices of $X.$ Then $Q$ has branch radius at most two.
Let $\Sigma$ be a surface without boundary and let $G$ be a graph drawn in $\Sigma$ without crossings. Towards simplifying notation, we do not distinguish the graph and its particular embedding, and we use $F(G)$ to denote the set of faces of $G,$ i.e., the set of connected components of the set $\Sigma\setminus G.$ Given $x,y\in V(G)\cup F(G),$ a radial $(x,y)$-path in $G$ of length at most $r$ is a sequence of vertices and faces of length $r+1,$ where faces and vertices alternate, which starts from $x$ and finishes in $y,$ and such that if $f\in F(G)$ and $v\in V(G)$ are consecutive in the sequence, then $v$ is a point of the boundary of $f.$ The radial distance between $x$ and $y,$ denoted by $\mathsf{rdist}_{G,\Sigma}(x,y)$ is the minimum length of a radial $(x,y)$-path in the embedding of $G$ in $\Sigma.$ The radial radius of $G$ is at most $r$ if there is an $x\in V(G)\cup F(G)$ such that for every $y\in V(G)\cup F(G),$ $\mathsf{rdist}_{G,\Sigma}(x,y)\leq r.$
The following has been proven in different forms, see eg, [10, 18, 24].
[…]
Let $G$ be a graph drawn in a surface $\Sigma$ of Euler-genus $g.$ There is a universal constant $c$ such that if the treewidth of G is more than $cr(g+1),$ then $G$ contains an $r$-wall $W$ as a subgraph which is a subset of a contractible closed disk of $\Sigma$ whose boundary is the perimeter of $W.$
[…]
Let $G$ be a graph drawn in a surface $\Sigma$ of Euler-genus at most $k$ with radial radius at most $r.$ Then ${\sf tw}(G)\in\mathcal{O}(g\cdot r)$
Proof.
Let $q$ be the maximum integer such that $G$ contains a $2(q+1)$-wall $W$ as a subgraph that is a subset of a closed disk of $\Sigma$ whose boundary is the perimeter of $W.$ Notice that the radial radius of $W$ is more than $q,$ which in turn implies that also the radial radius of $G$ is at least $q.$ As $G$ has radial radius at most $r,$ we have that $q\leq r.$ This, combined with Subsection 6.1 implies that ${\sf tw}(G)\in\mathcal{O}(g\cdot r).$
6.2 Bounding the biggest grid
Let us denote the $(k\times k)$-grid by $\Gamma_{k}.$ The perimeter of $\Gamma_{k},k\geq 3,$ is defined in the same way as it is defined for walls. We say that the $1$st layer of $\Gamma_{k}$ is its perimeter. Given some $s\leq\lfloor k/2\rfloor$ we define the $s$-th layer of $\Gamma_{k}$ as the perimeter of the $(k-2s+2\times k-2s+2)$-grid remaining if we remove the first $s-1$ layers of $\Gamma_{k}.$ We refer to the first $s$ layers of $\Gamma_{k}$ as its $s$ outermost layers.
Given an $(m\times m)$-grid $\Gamma_{m},m\geq 3,$ and some $k$ where $3\leq k\leq m$ and such that $m-k$ is even, we define the central $(k,k)$-subgrid of $\Gamma_{m}$ as the unique subgraph of $\Gamma_{m}$ obtained by removing its ${\frac{m-k}{2}}$ outermost layers.
Let $G$ and $H$ be graphs, $X\subseteq V(G)$, and assume that $H$ is an $X$-minor of $G,$ certified by the collection $\mathcal{S}=\{S_{v}\mid v\in V(H)\}$ of pairwise-disjoint connected subsets of $V(G).$ If $x\in S_{v},$ then we say that $x$ is a precursor of $v$ in $G$ and that $v$ is the heir of $x$ in $H.$
[…]
Let $(H,Y)$ and $(G,X)$ be annotated graphs where $H$ is an $X$-minor of $G$. Then ${\sf bg}(G,X)\geq{\sf bg}(H,Y).$
An important tool for the study of ${\sf bg}$ is the following.
[…]
Let $\Gamma_{m}$ be the $(m\times m)$-grid and let $S$ be a subset of the vertices in the central $(m-2\ell)\times(m-2\ell)$-subgrid of $\Gamma_{m},$ where =|$\ell=\lfloor\sqrt[4]{|S|}\rfloor.$ $G$ contains the $(\ell\times\ell)$-grid as an $S$-minor.
[…]
Let $G$ be a graph and let $X_{1},\ldots,X_{r}\subseteq V(G)$ such that, for $i\in[r],$ ${\sf bg}(G,X_{i})\leq\ell.$ Then ${\sf bg}(G,X_{1}\cup\cdots\cup X_{r})\in\mathcal{O}(r^{4}\ell^{2}).$
Proof.
Let $X=X_{1}\cup\cdots\cup X_{r}$ and let $q={\sf bg}(G,X).$ Let also $\Gamma_{q}$ be a $(q\times q)$-grid $X$-minor in $G,$ certified by a collection $\mathcal{S}=\{S_{v}\mid v\in V(\Gamma_{q})\}$ of pairwise-disjoint connected subsets of $V(G).$ We choose $i$ so that the cardinality of the set $Z_{i}\coloneqq\{u\in V(\Gamma_{q})\mid X_{i}\cap S_{u}\neq\emptyset\}$ is maximized, i.e., $X_{i}$ has the majority of intersections with the sets of $\mathcal{S}.$ As we fix $i,$ for simplicity we use $Z=Z_{i}.$
Clearly $Z$ is a subset of $V(\Gamma_{q})$ where $|Z|\geq q^{2}/r.$ For each (positive) margin $b,$ we define the set $B_{b}$ as the union of the $b$ outermost layers of $\Gamma_{q}.$ We now fix some $b\in[\lfloor q/2\rfloor],$ and observe that at most $4b(q-b)$ of the vertices of $\Gamma_{q}$ are in the outermost $b$ layers. This means that at least $f(q,b,r)\coloneqq\frac{q^{2}}{r}-4b(q-b)$ vertices of $Ζ$ are in the central $(q-2b)\times(q-2b)$-subgrid of $\Gamma_{m}.$
We choose the margin $b_{q,r}$ as the maximum integer $b$ such that $4]{f(q,b,r)}\rfloor\geq 2b.$
This implies that there are at least $f(q,b_{q,r},r)$ vertices from $Z$ in the $(q-2b_{q,r})\times(q-2b_{q,r})$-subgrid of $\Gamma_{q}.$ By Subsection 6.2 we obtain, ${\sf bg}(G,X_{i})\geq{\sf bg}(\Gamma_{q},Z_{i})\geq\lfloor\sqrt[4]{f(q,b_{q,r},r)}\rfloor\geq 2b_{q,r}\in\Omega(\frac{q^{1/2}}{{r}^{1/4}}).$ Therefore, ${\sf bg}(G,X_{1}\cup\cdots\cup X_{r})\in\mathcal{O}(r^{4}\ell^{2}).$
The next observation follows directly from the definition of ${\sf bg}.$
[…]
Let $G$ be a graph, $A\subseteq V(G),$ and $X\subseteq V(G)\setminus A,$ where ${\sf bg}(G-A,X)\leq k.$ Then ${\sf bg}(G,X\cup A)\leq k+|A|.$
[…]
For every graph $G,$ every $k\in\mathbb{N},$ and every surface $\Sigma$ with Euler-genus at most $g,$ if $A\subseteq V(G),$ $\delta$ is a $\Sigma$-decomposition of $G-A$ of width at most $w$ and breadth at most $b$, and ${\displaystyle X\coloneqq A\cup\bigcup\{\Sigma(c)\mid\mbox{\small$c$ is a vortex of $\delta$}\}},$ then ${\sf bg}(G,X)\in\mathcal{O}(b^{4}\cdot(b\cdot g\cdot w)^{2}+|A|).$
Proof.
Let $c_{1},\ldots,c_{b}$ be the vortices of $\delta.$ Also, let ${(\Sigma(c_{i}),\Omega_{i})}$ be the vortex society of $c_{i},$ $i\in[b].$ Because of Subsection 6.2, we may assume that $A=\emptyset$ and prove that ${\sf bg}(G,X)\in\mathcal{O}(b^{4}\cdot(b\cdot g\cdot w)^{2}).$ Because of Subsection 6.2, it is enough to prove that for every $\hat{c}\in\{c_{1},\ldots,c_{b}\},$ if $S=V(\Sigma(\hat{c})),$ then ${\sf bg}(G,S)\in\mathcal{O}(bgw).$
Let $z=\lfloor{\sf bg}(G,S)/{c^{\prime}}\rfloor$ where $c^{\prime}$ is the constant of Subsection 6.1.
According to Subsection 6.1 $G$ contains a hairy annotated $z$-wall $(\hat{W},S).$ Moreover, ${\sf tw}(\hat{W})\geq z.$
As a first step, we enhance $G$ by adding, for each vortex $c_{i}\in\{c_{1},\ldots,c_{b}\},$ a cycle $C_{i}$ joining every pair of consecutive vertices in $\Omega_{i}.$ We further enhance $G$ by adding a new vertex $a,$ called apex vertex, and make it adjacent to all the vertices of $S.$ We also update $\delta$ by setting $\Sigma(c_{i})\coloneqq\Sigma(c_{i})\cup C_{i},$ for every $i\in[b],$ and we further enhance $\Sigma(\hat{c})$ by adding $a$ and making it adjacent to all vertices of $S.$
Notice that the above enhancements cannot increase the width of the vortices of $\delta$ by more than three. We now define the pair $(W^{+},S)$ where $W^{+}$ is obtained if we first consider the graph ${W}\cup\boldsymbol{\boldsymbol{\bigcup}}\{C_{i}\mid i\in[b]\}$ and then add the vertex $a$ and all the edges between $a$ and $S.$ Clearly, $W^{+}$ may have vertices outside $\hat{c},$ however all vertices in $\{a\}\cup S$ are vertices of $\Sigma(\hat{c}).$ We now consider the graphs
  • $H_{1}=W^{+}\setminus\bigcup\{\Sigma(c_{i})\setminus V(C_{i})\mid i\in[b]\},$ and
  • $H_{2}=H_{1}\cup\bigcup\{\Sigma(c_{i})\mid i\in[b]\}.$
Observe that $H_{1}$ is a subgraph of $W^{+}$ and $W^{+}$ is a subgraph of $H_{2}.$ By construction, $H_{1}$ is a graph that is embedded in $\Sigma$ without crossings. Also keep in mind that ${\sf tw}(W^{+})\geq{\sf tw}(W)\geq z.$
Recall that every branch vertex of $W$ is at branch distance at most three from $a$. This fact combined with the construction of $H_{1}$ implies that every face or vertex of $H_{1}$ is at radial distance $\mathcal{O}(b)$ from the face of $H_{1}$ whose boundary of $C_{\hat{c}}.$ We obtain that the radius of $H_{1}$ is $\mathcal{O}(b)$ and this, combined with Subsection 6.1 implies that ${\sf tw}(H_{1})\in\mathcal{O}(bg).$
Our next step is to use the fact that ${\sf tw}(H_{1})\in\mathcal{O}(bg)$ in order to prove that ${\sf tw}(H_{2})\in\mathcal{O}(bgw).$ For this, let $(T,\beta)$ be a tree-decomposition of $H_{1}$ and let $L^{i}=\langle X_{1}^{i},X_{2}^{i},\dots,X_{n_{i}}^{i},x_{1}^{i},x_{2}^{i},\dots,x_{n_{i}}^{i}\rangle$ be a linear decomposition of $\Sigma(c_{i}),$ assuming that $x_{1}^{i},x_{2}^{i},\dots,x_{n_{i}}^{i}$ are the vertices of $\Omega_{i}$ ordered according to $\Omega_{i}.$ We now build a tree-decomposition $(T,\beta^{\prime})$ of $H_{2}$ where for each $t\in V(T),$ we define $\beta^{\prime}(t)$ to be the union of the set $\beta(t)$ and every set $X_{j}^{i},$ for which $x_{j}^{i}$ is a vertex of $\beta(t)\cap\boldsymbol{\boldsymbol{\bigcup}}\{\{x_{1}^{i},x_{2}^{i},\dots,x_{n_{i}}^{i}\}\mid i\in[b]\}.$ As each $X_{j}^{i}$ has $\mathcal{O}(w)$ vertices, it follows that $(T,\beta^{\prime})$ has width $\mathcal{O}(bgw),$ as required. We conclude that ${\sf tw}(H_{2})\in\mathcal{O}(bgw).$
Finally, recall that $W^{+}$ is a subgraph of $H_{2},$ therefore $z\leq{\sf tw}(W^{+})\leq{\sf tw}(H_{2})\in\mathcal{O}(bgw).$ In other words, ${\sf bg}(G,S)=c^{\prime}z^{2}\in\mathcal{O}(bgw)$ as required. ∎
Let $α$ be an integer such that for every graph $G,$ ${\sf tw}(G)\in\mathcal{O}({\sf bg}(G))^{\alpha}.$ Such a constant $α$ exists by Subsection 3.1. Then Subsection 6.2 implies the following.
[…]
For every graph $G,$ every $k\in\mathbb{N},$ and every surface $\Sigma$ with Euler-genus at most $g,$ if $A\subseteq V(G),$ $\delta$ is a $\Sigma$-decomposition of $G-A$ with width $\leq w$ and breadth $\leq b$ and ${\displaystyle X=A\cup\bigcup\{\Sigma(c)\mid\mbox{\small$c$ is a vortex of $\delta$}\}},$ then ${\sf tw}(G,X)\in\mathcal{O}((b^{4}\cdot(b\cdot g\cdot w)^{2}+|A|)^{α}),$ where $α$ is the constant of Subsection 3.1.
7 The lower bound
Using the concept of clique-sum closures, we can easily generate new graph parameters. For this, given a graph parameter ${\sf p}\colon\mathcal{G}_{\text{all}}\to\mathbb{N}$ we define the graph parameter ${\sf p}^{*}$ where for every graph $G,$ ${\sf p}^{*}(G)$ is the minimum $k$ such that $G$ is in the clique-sum closure of the class of graphs where the value of ${\sf p}$ is at most $k.$ Notice that ${\sf p}^{*}=({\sf p}^{*})^{*}.$ It is easy to see that if size is the parameter returning the number of vertices of a graph minus one, then ${\sf tw}=\textsf{size}^{*}.$
[…]
If ${\sf p}$ is a minor-monotone graph parameter, then ${\sf p}^{*}$ is also minor-monotone.
Given a graph parameter ${\sf p}\colon\mathcal{G}_{\text{all}}\to\mathbb{N}$ and a minor-closed graph class $\mathcal{G}$ we define the modulator composition of ${\sf p}$ and $\mathcal{G}$, as the parameter ${\sf p}_{\mathcal{G}}:\mathcal{G}_{\text{all}}\to\mathbb{N}$, so that
${\sf p}_{\mathcal{G}}(G)$$\coloneqq$ $\min\{k\mid\mbox{$\mbox{there exists an $X\subseteq V(G)$ such that\leavevmode\nobreak\ }{\sf p}(G,X)\leq k\mbox{\leavevmode\nobreak\ and\leavevmode\nobreak\ }G-X\in\mathcal{G}$}\}.$ (12)
Notice that, given that ${\sf p}$ is minor-monotone and $\mathcal{G}$ is minor closed, we have that ${\sf p}_{\mathcal{G}}$ is minor-monotone as well. Using (12), we define the “annotated modulation to a surface”-version of ${\sf p}$ as the graph parameter ${\sf p}_{\Sigma}$ where ${\sf p}_{\Sigma}\coloneqq{\sf p}_{\mathcal{E}_{\Sigma}}$. Notice that, because of Subsection 3.1, ${\sf tw}\sim_{\mathsf{P}}{\sf bg}.$ By the above definitions, one can easily deduce that ${\sf tw}_{\Sigma}\sim_{\mathsf{P}}{\sf bg}_{\Sigma}$ and moreover that
${\sf tw}_{\Sigma}^{*}$ $\sim_{\mathsf{P}}$ ${\sf bg}_{\Sigma}^{*}$ (13)
In (13), we agree that ${\sf tw}_{\Sigma}^{*}$ is a shortcut of $({\sf tw}_{\Sigma})^{*}$ and that ${\sf bg}_{\Sigma}^{*}$ is a shortcut of $({\sf bg}_{\Sigma})^{*}$. The equivalence in (13) will permit us to use ${\sf bg}_{\Sigma}$ instead of ${\sf tw}_{\Sigma},$ to prove our lower bounds.
In this section we consider for fixed values of $h$ and $c,$ the Dyck-grid $\mathscr{D}_{k}^{(h,c)}$ and a surface $\Sigma$ where $\mathscr{D}_{k}^{(h,c)}$ cannot be embedded (that is a surface $\Sigma$ not containing $\Sigma^{(h,c)}$). Our target is to prove that $\mathscr{D}_{k}^{(h,c)}$ cannot be the clique-sum closure of graphs with low-treewidth modulators to $\Sigma.$ In other words, we prove a (polynomial) lower bound to ${\sf tw}_{\Sigma}^{*}(\mathscr{D}_{k}^{(h,c)})$ or, alternatively, because of (13), a lower bound to ${\sf bg}_{\Sigma}^{*}(\mathscr{D}_{k}^{(h,c)})$ (in terms of $k$).
Our proof has two steps. In Subsection 7.1 we prove that ${\sf bg}_{\Sigma}(\mathscr{D}_{k}^{(h,c)})\in\Omega(k^{1/5}).$ Based on this result, in Subsection 7.2 we prove that ${\sf bg}_{\Sigma}^{*}(\mathscr{D}_{k}^{(h,c)})\in\Omega(k^{1/480}).$ In our proofs we do not make any particular effort to optimize the parametric dependences. While we keep them being polynomial, it seems that much better polynomial bounds might be achieved with a more fine-grained analysis.
7.1 Modulators of Dyck-grids
In our proof we work with a simpler version of $\mathscr{D}^{(h,c)},$ where the annulus grid is suppressed. The new parametric graph is denoted by $\tilde{\mathscr{D}}^{(h,c)}=\langle\tilde{\mathscr{D}}_{k}^{(h,c)}\rangle_{k\in\mathbb{N}},$ where the graph $\tilde{\mathscr{D}}_{k}^{(h,c)}$ is obtained from $\mathscr{D}_{k}^{(h,c)}$ by removing the set $S$ of $4k^{2}$ vertices corresponding to the annulus grid $\mathscr{A}_{k}$ and, in the resulting graph, among the $2k$ neighbors of the vertices that have been removed, we join by edges the $k$ pairs that belong to the same cycle of $\mathscr{D}_{k}^{(h,c)}$ (see Figure 2 for an example). Clearly $\tilde{\mathscr{D}}_{k}^{(h,c)}\leq\mathscr{D}_{k}^{(h,c)}$ and it is easy to see that ${\mathscr{D}}_{k}^{(h,c)}\leq\tilde{\mathscr{D}}_{2k}^{(h,c)}.$ This implies the following observation.
[…]
For every $(h,c)\in\mathbb{N}\times[0,2]\setminus\{(0,0)\},$ $\mathscr{D}^{(h,c)}\sim_{{\mathsf{L}}}\tilde{\mathscr{D}}^{(h,c)}.$
[…]
Figure 14: The graph $\tilde{\mathscr{D}}_{8}^{1,2}$ and a half-integral packing of four subdivisions of $\tilde{\mathscr{D}}_{2}^{1,2},$ colored by different colors.
[…]

### v4

@@ 7 From local to global structure theorems
We prove that the following, slightly stronger, version of the first claim holds by induction on the number of vertices in $G.$
Strengthened claim: Let $G$ be a graph which does not have a minor isomorphic to $\mathscr{H}^{(j)}_{t_{j}}$ for any $j\in[k]$ and let $Z\subseteq V(G)$ be a set of size at most $3\alpha.$ Then $G$ has a tree decomposition $(T,\beta)$ with root $f,$ where $Z\subseteq\beta(f),$ and which is of adhesion less than $\mathbf{a}(g_{1},g_{2},k,r)$ such that for every $d\in V(T)$ either
  • 1.
    $|\beta(d)|\leq 4\mathbf{a}(g_{1},g_{2},k,r),$ or
  • 2.
    there exists a set $A_{d}\subseteq\beta(d)$ of size at most $4\mathbf{a}(g_{1},g_{2},k,r)$ and integers $\mathsf{h}_{d},\mathsf{c}_{d},$ $\mathsf{c}_{d}\in[0,2],$ together with a surface $\Sigma=\Sigma^{(\mathsf{h}_{d},\mathsf{c}_{d})}$ where $\mathscr{D}^{(\mathsf{h}_{1},0)}_{k}$ does not embed and, if $\mathsf{c}_{2}\neq 0,$ where also $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k}$ does not embed such that for the torso $G_{d}$ of $G_{d}$ at $d,$ the graph $G-A_{d}$ has a $\Sigma$-decomposition $\delta$ of width at most $\mathbf{a}(g_{1},g_{2},k,r)$ and breadth at most $\mathbf{v}_{\mathsf{b},g_{1},g_{2}}$ such that, if $X_{d}$ is the set of all vertices of $G_{d}-A_{d}$ drawn in the interior of the vortices of $\delta$ together with the vertices in $A_{d},$
    1. 1.
      there exists an $r$-wall $W_{d}$ which is flat in $\delta,$
    2. 2.
      there is no vertex of $G_{d}-A_{d}$ which is drawn in the interior of a non-vortex cell of $\delta,$
    3. 3.
      $G_{d}-X_{d}$ is a minor of $G,$ and
    4. 4.
      for every neighbour $d^{\prime}$ of $d$ in $T$ it holds that $|\big{(}\beta(d)\cap\beta(d^{\prime})\big{)}\setminus\big{(}A_{d}\cup X_{d}\big{)}|\leq 3.$
Assume towards a contradiction that the strengthened claim is false. We choose $G,$ $Z,$ and $k,\mathsf{h}_{1},\mathsf{h}_{2},\mathsf{c}_{2},r$ to form a counterexample minimizing $|V(G)|+|V(G)\setminus Z|.$ It follows that $|V(G)|>4\mathbf{a}(g_{1},g_{2},k,r)$ since otherwise we could choose $(T,\beta)$ to be a trivial tree decomposition with a single node and a unique bag containing all of $V(G).$ Moreover, we have $|Z|=3\mathbf{a}(g_{1},g_{2},k,r)$ since otherwise we could add an arbitrary vertex of $V(G)\setminus Z$ to $Z$ and thereby contradict our choice for the counterexample.
Suppose there is a set $X\subseteq V(G)$ of order less than $\mathbf{a}(g_{1},g_{2},k,r)$ such that $|V(J)\cap Z|\leq\lfloor\frac{2}{3}|Z|\rfloor\leq\mathbf{a}(g_{1},g_{2},k,r).$ In this case let us introduce a node $f$ and set $\beta(f)\coloneqq X\cup Z.$ Now, for each component $J$ of $G-X$ let $G_{J}\coloneqq G[V(J)\cup X]$ and $Z_{J}^{\prime}\coloneqq Z\cap V(J).$ Notice that $|Z_{J}^{\prime}|<3\mathbf{a}(g_{1},g_{2},k,r).$ In case $V(G_{J})=Z_{J}^{\prime}$ there is nothing to show here, otherwise we may pick an arbitrary vertex $v_{J}\in V(G_{J})\setminus Z_{J}^{\prime}$ and set $Z_{J}\coloneqq Z_{J}^{\prime}\cup\{v_{J}\}.$ Observe that $G_{J},$ $Z_{J},$ and $k,\mathsf{h}_{1},\mathsf{h}_{2},\mathsf{c}_{2},r$ do not form a counterexample by our minimality assumption. Hence, there exists a tree decomposition $(T_{i},\beta_{i})$ with root $f_{i}$ and $Z_{J}\subseteq\beta(f_{i})$ which satisfies all requirements of the strengthened claim. We may form a tree decomposition $(T,\beta)$ by joining $f$ to the vertices $f_{1}$ and $f_{2},$ thereby creating the tree $T.$ It is straight forward to check that $(T,\beta)$ meets all requirements of the strengthened claim which is a direct contradiction to the choice of our counterexample. Observe that it is possible to determine if such a decomposition exists, at least with a slightly worse approximation factor, by using the techniques from [2].
[…]